OpenStudy (anonymous):

A newly discovered solar system is found 150 billion kilometers from the sun. How many AU does this represent? Show your work

OpenStudy (festinger):

From google: 1 Astronomical Unit =149 597 871 kilometers Cancel "units" by arraging the above in a way that kilometers get cancelled off, leaving AU. \[\frac{1AU}{149597871km}=1\] \[150*10^{9}km=150*10^{9}km*1=150*10^{9}km*\frac{1AU}{149597871km}=1003AU\]